# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

### v4

@@ Abstract
A long-standing belief holds that Bayesian Optimization (BO) with standard Gaussian processes (GP) — referred to as standard BO — underperforms in high-dimensional optimization problems. While this belief seems plausible, it lacks both robust empirical evidence and theoretical justification. To address this gap, we present a systematic investigation. First, through a comprehensive evaluation across eleven widely used benchmarks, we found that while the popular Square Exponential (SE) kernel often leads to poor performance, using Matérn kernels enables standard BO to consistently achieve top-tier results, frequently surpassing methods specifically designed for high-dimensional optimization. Second, our theoretical analysis reveals that the SE kernel’s failure primarily stems from improper initialization of the length-scale parameters, which are commonly used in practice but can cause gradient vanishing in training. We provide a probabilistic bound to characterize this issue, showing that Matérn kernels are less susceptible and can robustly handle much higher dimensions. Third, we propose a simple robust initialization strategy that dramatically improves the performance of the SE kernel, bringing it close to state-of-the-art methods, without requiring any additional priors or regularization. We prove another probabilistic bound that demonstrates how the gradient vanishing issue can be effectively mitigated with our method. Our findings advocate for a re-evaluation of standard BO’s potential in high-dimensional settings.
@@ 1 Introduction
Many applications require optimizing complex functions, allowing queries only for the function values, possibly with noises, yet without any gradient information. Bayesian Optimization (BO) [27, 21] is a popular approach for addressing such challenges. BO typically employs Gaussian process (GP) [32] as a probabilistic surrogate. It iteratively approximates the target function, integrates the posterior information to maximize an acquisition function for generating new inputs at which to query, then updates the GP model with new examples, and concurrently approach the optimum.
Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems [8, 23, 6, 22, 18, 31, 19]. It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure [8, 23]. A continuous line of research is dedicated to developing novel high-dimensional BO methods. The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs. One class of methods assumes a decomposition structure within the functional space [16, 24, 10, 33], where the target function is expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables. Another family of methods assumes that the inputs are intrinsically low-rank. These methods either project the original inputs into a low-dimensional space [31, 23, 18] or use sparse-inducing priors to trim down massive input variables [6]. Subsequently, GP surrogates are built with the reduced input dimensions.
@@
  • Empirical Results. We investigated BO with standard GP across eleven widely-used benchmarks, including six synthetic and five real-world high-dimensional optimization tasks. The number of variables ranged from 30 to 388. We compared standard BO with nine state-of-the-art high-dimensional BO methods, and performed extensive evaluations and ablation studies. Surprisingly, while the popular SE kernel often led to poor performance, switching to the ARD Matérn kernel enabled standard BO to consistently achieve the best (or near-best) optimization performance across all the tasks. When successful, standard BO seems more flexible in accommodating various structures within the target functions.
@@ 2 Standard Bayesian Optimization
Consider maximizing a $d$-dimensional objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$. The form of $f(\cdot)$ is unknown. We can query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, but no gradient information is accessible. We aim to find ${\bf x}^{\dagger}=\mathrm{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$. To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to predict $f$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction. This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate. The utility typically encapsulates an exploration-exploitation trade-off. By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum. Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy. BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model. Then the iterative procedure repeats until convergence or a predefined stopping criterion is met.
The standard BO adopts GP regression [32] for surrogate modeling. A GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is usually chosen as a Mercer kernel function. One most popular kernel used in BO is the Square Exponential (SE) kernel with Automatic Relevance Determination (ARD),
@@
Given training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$, let us denote the function values at ${\bf X}$ as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, which according to the GP prior, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Then one can employ a Gaussian likelihood for the observations ${\bf y}$, and obtain
@@
where $\sigma^{2}$ is the noise variance. To estimate the GP parameters, e.g., the length-scales $\bm{\ell}$, one can maximize the marginal likelihood or use MCMC sampling. The predictive distribution (given the GP parameters) at any new input ${\bf x}^{*}$, is a closed-form, conditional Gaussian distribution, $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right)$, where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
In each iteration, given the current GP surrogate model, an acquisition function is maximized to identify the next input location at which to query. One commonly used acquisition function is the upper confidence bound (UCB), defined as $\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X}$, where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) [15], Thompson sampling (TS) [25], the recently proposed log-EI [1], among others.
@@ 3 High Dimensional Bayesian Optimization
where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$. There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$. In [16], the input ${\bf x}$ is partitioned into non-overlapped groups. After every a few BO iterations, a new partition is selected from a set of random partitions. The selection is based on the model evidence. In [24], the variable groups can overlap, and are represented as the maximum cliques on a dependency graph. A Gibbs sampling method was developed to learn the structure of the dependency graph. Han et al., [10] proposed Tree-UCB, which restricted the dependency graph to be tree-structured so as to boost the efficiency of structure learning. However, the more recent work [33] points out that, learning the group structure through relatively small data can be misleading. A wrongly learned structure can cause queries to be stuck at local optima. The work proposes RDUCB that randomly decomposes ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
Structural Assumption in Input Space. Another class of methods assumes a low-rank structure in the input space. Many of them introduce Low-dimensional Embeddings. Wang et al., [31] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al., [23] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially imposes further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  [18]. ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue. Another direction is to impose Partial Variable Dependency by triming down the input variables. The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [12]. While one can also use MAP estimation, the performance is often much worse.
@@ 4 Theoretical Analysis
To understand why standard BO can fail in high-dimensional cases, in particular with the SE kernel (see Section 6), we look into the gradient structure during the GP training. Specifically, we consider at the beginning, all the length-scale parameters are set to the same initial value, $\ell_{1}=\ldots=\ell_{d}=\ell_{0}$. From the marginal likelihood (3), we can derive the gradient w.r.t each length-scale,
@@
$\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ij}=\frac{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})}{\partial\ell_{k}}=\frac{2a}{\ell_{k}}\exp(-\frac{\|{\bf x}_{i}-{\bf x}_{j}\|^{2}}{\ell_{0}^{2}})\frac{(x_{ik}-x_{jk})^{2}}{\ell_{k}^{2}}\leq\frac{2a}{\ell_{k}}\cdot\frac{\rho^{2}}{e^{\rho^{2}}},$ (6)
@@
Intuitively, as the input dimension $d$ increases, the squared distance between input vectors grows rapidly. Consequently, the factor $\frac{\rho^{2}}{e^{\rho^{2}}}$ in (6) can quickly fall below the machine epsilon $\xi$, causing the gradient vanishing issue [3, 11]. That is, the length-scale cannot be effectively updated according to $\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}$, due to the limitations of the numerical precision.
@@
The lower bound grows exponentially with the input dimension $d$ and converges to one. This implies that given any fixed choice of $\ell_{0}$ relevant to $d$ (e.g., $\ell_{0}=1$), as $d$ increases, the probability $p(\rho\geq\tau_{\text{SE}})$ will rapidly approach one, leading to vanishing gradient for every $\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})$. As a consequence, each length scale $\ell_{k}$ in (5) cannot be effectively updated, preventing meaningful training [^1] A union bound can be directly derived from (7) to lower bound the joint probability that all gradients $\{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}\}_{i,j}$ across $N$ training instances are below machine epsilon. However, this bound is too loose to reflect our practical observation (see Fig. 1 and  2 and Table 2). Therefore, we focus our discussion on a single pair of inputs. . The inferior prediction accuracy of the GP model further contributes to poor performance of the BO procedure. Since $\tau_{\text{SE}}$ is small, gradient vanishing can occur even when $d$ is not large. For example, for $\ell_{0}=0.5$ and $\ell_{0}=1$, we have $p(\rho>\tau_{\text{SE}})>0.99$ when $d\geq 205$ and $d\geq 473$, respectively.
@@
Numerical Verification. To verify whether gradient vanishing occurs during practical GP training and to access whether our analysis aligns with the empirical observations, we conducted numerical experiments using two synthetic functions, Hartmann6($d$, 6) and Rosenbrock(d,d) (see Section 6.1 for definitions). We varied the input dimension $d$ across {50, 100, 200, 300, 400, 500, 600}, with the domain set to $[0,1]^{d}$. For each value of $d$, we uniformly sampled $500$ training inputs and $100$ test inputs. We trained GP with both SE and Matérn kernels, using length-scale initialization $\ell_{0}$ from $\{0.1,0.5,0.693,1.0,\sqrt{d}\}$. Note $\ell_{0}=0.693=\text{SoftPlus}(0)$ is a popular initialization choice, used as the default choice in influential GP/BO libraries GPyTorch [9] and BOTorch [2]. Training and testing were repeated 20 times for each $d$ and $\ell_{0}$. We then evaluated (1) the average relative $L_{2}$ difference between the length-scale vector before and after training, namely, $\|\bm{\ell}_{\text{trained}}-\bm{\ell}_{\text{init}}\|/\|\bm{\ell}_{\text{init}}\|$, (2) the average gradient norm at the first training step, and (3) the average test Mean-Squared-Error (MSE).

### v5

# Standard Gaussian Process is All You Need for High-Dimensional Bayesian Optimization

@@ Abstract
A long-standing belief holds that Bayesian Optimization (BO) with standard Gaussian processes (GP) — referred to as standard BO — underperforms in high-dimensional optimization problems. While this belief seems plausible, it lacks both robust empirical evidence and theoretical justification. To address this gap, we present a systematic investigation. First, through a comprehensive evaluation across twelve benchmarks, we found that while the popular Square Exponential (SE) kernel often leads to poor performance, using Matérn kernels enables standard BO to consistently achieve top-tier results, frequently surpassing methods specifically designed for high-dimensional optimization. Second, our theoretical analysis reveals that the SE kernel’s failure primarily stems from improper initialization of the length-scale parameters, which are commonly used in practice but can cause gradient vanishing in training. We provide a probabilistic bound to characterize this issue, showing that Matérn kernels are less susceptible and can robustly handle much higher dimensions. Third, we propose a simple robust initialization strategy that dramatically improves the performance of the SE kernel, bringing it close to state-of-the-art methods, without requiring additional priors or regularization. We prove another probabilistic bound that demonstrates how the gradient vanishing issue can be effectively mitigated with our method. Our findings advocate for a re-evaluation of standard BO’s potential in high-dimensional settings. The code is released at https://github.com/XZT008/Standard-GP-is-all-you-need-for-HDBO.
@@ 1 Introduction
Many applications require optimizing complex functions, allowing queries only for the function values, possibly with noises, yet without any gradient information. Bayesian Optimization (BO) (Snoek et al., 2012; Mockus, 2012) is a popular approach for addressing such challenges. BO typically employs Gaussian process (GP) (Williams & Rasmussen, 2006) as a probabilistic surrogate. It iteratively approximates the target function, integrates the posterior information to maximize an acquisition function for generating new inputs at which to query, then updates the GP model with new examples, and concurrently approach the optimum.
Despite numerous successes, there has been a widespread belief that BO with standard GP regression, referred to as standard BO, is limited to low-dimensional optimization problems (Frazier, 2018; Nayebi et al., 2019; Eriksson & Jankowiak, 2021; Moriconi et al., 2020; Letham et al., 2020; Wang et al., 2016; Li et al., 2016). It is commonly thought that the number of optimization variables should not exceed 15 or 20, as beyond this threshold, BO is prone to failure (Frazier, 2018; Nayebi et al., 2019). A continuous line of research is dedicated to developing novel high-dimensional BO methods. The fundamental strategy is to impose strong structural assumptions into surrogate modeling so as to avoid directly dealing with high-dimensional inputs. One class of methods assumes a decomposition structure within the functional space (Kandasamy et al., 2015; Rolland et al., 2018; Han et al., 2021; Ziomek & Ammar, 2023), where the target function is expressed as the summation of a group of low-dimensional functions, each operating over a small number of variables. Another family of methods assumes that the inputs are intrinsically low-rank. These methods either project the original inputs into a low-dimensional space (Wang et al., 2016; Nayebi et al., 2019; Letham et al., 2020) or use sparse-inducing priors to trim down massive input variables (Eriksson & Jankowiak, 2021). Subsequently, GP surrogates are built with the reduced input dimensions.
@@
  • Empirical Results. We investigated BO with standard GP across eleven widely used benchmarks and one novel benchmark, encompassing six synthetic and six real-world high-dimensional optimization tasks. The number of variables ranged from 30 to 1,003. We compared standard BO with nine state-of-the-art high-dimensional BO methods, and performed extensive evaluations. Surprisingly, while the popular SE kernel often led to poor performance, switching to the ARD Matérn kernel enabled standard BO to nearly always achieve the best (or near-best) optimization performance. When successful, standard BO seems more flexible in accommodating various structures within the target functions.
@@ 2 Standard Bayesian Optimization
Consider maximizing a $d$-dimensional objective function $f:\mathcal{X}\subset\mathbb{R}^{d}\rightarrow\mathbb{R}$, where the function form is unknown. We can query the function value at any input ${\bf x}\in\mathcal{X}$, possibly with noises, but no gradient information is accessible. We aim to find ${\bf x}^{\dagger}=\mathrm{argmax}_{{\bf x}\in\mathcal{X}}f({\bf x})$. To achieve this, Bayesian Optimization (BO) employs a probabilistic surrogate model to predict $f$ across the domain $\mathcal{X}$, while also quantifying the uncertainty of the prediction. This information is then integrated to compute an acquisition function, which measures the utility of querying at any new input location given the current function estimate. By maximizing the acquisition function, BO identifies a new input location at which to query, ideally closer to the optimum. Concurrently, the acquired new example is incorporated into the training dataset, and the surrogate model is retrained to improve the accuracy. BO begins by querying at a few (typically randomly selected) input locations, and trains an initial surrogate model. The iterative procedure repeats until convergence or a stopping criterion is met.
The standard BO adopts GP regression (Williams & Rasmussen, 2006) for surrogate modeling. A GP prior is placed over the target function, $f\sim\mathcal{GP}\left(m({\bf x}),\kappa({\bf x},{\bf x}^{\prime})\right)$, where $m({\bf x})$ is the mean function, which is often set as a constant, and $\kappa(\cdot,\cdot)$ is the covariance function, which is usually chosen as a Mercer kernel function. One most popular kernel used in BO is the Square Exponential (SE) kernel with Automatic Relevance Determination (ARD),
@@
Given training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$, let us denote the function values at ${\bf X}$ as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, which according to the GP prior, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Then one can employ a Gaussian likelihood for the observations ${\bf y}$, and
@@
where $\sigma^{2}$ is the noise variance. To estimate the GP parameters, e.g., the length-scales $\bm{\ell}$, one can maximize the marginal likelihood (3). The predictive distribution at any new input ${\bf x}^{*}$, is conditional Gaussian, $p\left(f({\bf x}^{*})|{\bf y}\right)=\mathcal{N}\left(f({\bf x}^{*})|\mu({\bf x}^{*}),v({\bf x}^{*})\right)$, where $\mu({\bf x}^{*})=m({\bf x}^{*})+\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$, and $v({\bf x}^{*})=\kappa({\bf x}^{*},{\bf x}^{*})-\kappa({\bf x}^{*},{\bf X})({\bf K}+\sigma^{2}{\bf I})^{-1}\kappa({\bf X},{\bf x}^{*})$.
In each iteration, given the current GP surrogate model, an acquisition function is maximized to identify the next input location at which to query. One commonly used acquisition function is the upper confidence bound (UCB), defined as $\text{UCB}({\bf x})=\mu({\bf x})+\lambda\sqrt{v({\bf x})},\;\;{\bf x}\in\mathcal{X}$, where $\lambda$ represents the exploration level. There have been other popular acquisition functions, such as Expected Improvement (EI) (Jones et al., 1998), Thompson sampling (TS) (Russo et al., 2018), the recently proposed log-EI (Ament et al., 2024), among others.
@@ 3 High Dimensional Bayesian Optimization
where each ${\bf x}^{j}\subset{\bf x}$ is a small group of input variables, and ${\bf x}={\bf x}^{1}\cup\ldots\cup{\bf x}^{M}$. There can be a variety of choices for the group number $M$ and each variable group ${\bf x}^{j}$. In (Kandasamy et al., 2015), the input ${\bf x}$ is partitioned into non-overlapped groups. After every a few BO iterations, a new partition is selected from a set of random partitions. The selection is based on the model evidence. In (Rolland et al., 2018), the variable groups can overlap, and are represented as the maximum cliques on a dependency graph. A Gibbs sampling method was developed to learn the structure of the dependency graph. Han et al. (2021) proposed Tree-UCB, which restricts the dependency graph to be tree-structured so as to boost the efficiency of structure learning. However, the more recent work (Ziomek & Ammar, 2023) points out that, learning the group structure through relatively small data can be misleading. A wrongly learned structure can cause queries to be stuck at local optima. This work proposes RDUCB that randomly decomposes ${\bf x}$ into a tree structure, and surprisingly works better than those learned structures through various methods.
Structural Assumption in Input Space. Another class of methods assumes a low-rank structure in the input space. Many of them introduce Low-dimensional Embeddings. Wang et al. (2016) proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al. (2019) proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially imposes further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  (Letham et al., 2020). ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue. Another direction is to impose Partial Variable Dependency by triming down the input variables. The recent start-of-the-art, SaasBO (Eriksson & Jankowiak, 2021), uses a horse-shoe prior (Carvalho et al., 2009) over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension.
@@ 4 Theoretical Analysis
To understand why standard BO can fail in high-dimensional cases, in particular with the SE kernel (see Section 6), we look into the gradient structure during the GP training. Specifically, we consider at the beginning, all the length-scale parameters are set to the same initial value, $\ell_{1}=\ldots=\ell_{d}=\ell_{0}$. From the marginal likelihood (3), we derive the gradient w.r.t each length-scale,
@@
$\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ij}=\frac{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})}{\partial\ell_{k}}=\frac{2a}{\ell_{k}}\exp(-\frac{\|{\bf x}_{i}-{\bf x}_{j}\|^{2}}{\ell_{k}^{2}})\frac{(x_{ik}-x_{jk})^{2}}{\ell_{k}^{2}}\leq\frac{2a}{\ell_{k}}\cdot\frac{\rho^{2}}{e^{\rho^{2}}},$ (6)
@@
Intuitively, as the input dimension $d$ increases, the squared distance between input vectors grows rapidly. Consequently, the factor $\frac{\rho^{2}}{e^{\rho^{2}}}$ in (6) can quickly fall below the machine epsilon $\xi$, causing the gradient vanishing issue (Bengio et al., 1994; Hanin, 2018). That is, the length-scale cannot be effectively updated according to $\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}$, due to the limitations of the numerical precision.
@@
The lower bound grows exponentially with the input dimension $d$ and converges to one. This implies that given any fixed choice of $\ell_{0}$ relevant to $d$ (e.g., $\ell_{0}=1$), as $d$ increases, the probability $p(\rho\geq\tau_{\text{SE}})$ will rapidly approach one, leading to vanishing gradient for every $\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})$. As a consequence, each length scale $\ell_{k}$ in (5) cannot be effectively updated, preventing meaningful training [^1] A union bound can be directly derived from (7) to lower bound the joint probability that all gradients $\{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}\}_{i,j}$ across $N$ training instances are below machine epsilon. However, this bound is too loose to reflect our practical observation (see Fig. 1 and  2 and Table 2). Therefore, we focus our discussion on a single pair of inputs. . The inferior prediction accuracy of the GP model further contributes to poor performance of the BO procedure. Since $\tau_{\text{SE}}$ is small, gradient vanishing can occur before $d$ becomes extremely large. For example, for $\ell_{0}=0.5$ and $\ell_{0}=1$, the probability $p(\rho>\tau_{\text{SE}})$ exceeds $0.99$ when $d\geq 205$ and $d\geq 473$, respectively.
@@
Numerical Verification. To verify whether gradient vanishing occurs during practical GP training and to access whether our analysis aligns with the empirical observations, we conducted numerical experiments using two synthetic functions, Hartmann6($d$, 6) and Rosenbrock(d,d) (see Section 6.1 for definitions). We varied the input dimension $d$ across {50, 100, 200, 300, 400, 500, 600}, with the domain set to $[0,1]^{d}$. For each value of $d$, we uniformly sampled $500$ training inputs and $100$ test inputs. We trained GP with both SE and Matérn kernels, using length-scale initialization $\ell_{0}$ from $\{0.1,0.5,0.693,1.0,\sqrt{d}\}$. Note $\ell_{0}=0.693=\text{SoftPlus}(0)$ is a popular initialization choice, used as the default choice in influential GP/BO libraries GPyTorch (Gardner et al., 2018) and BOTorch (Balandat et al., 2020). Training and testing were repeated 20 times for each $d$ and $\ell_{0}$. We then evaluated (A) the average relative $L_{2}$ difference between the length-scale vector before and after training, namely, $\|\bm{\ell}_{\text{trained}}-\bm{\ell}_{\text{init}}\|/\|\bm{\ell}_{\text{init}}\|$, (B) the average gradient norm at the first training step, and (C) the average test Mean-Squared-Error (MSE).